OpenStudy (anonymous):

Insert five geo metric means between 1 and 1000. Please show working !

OpenStudy (anonymous):

1,a,b,c,d,e,1000 let these variables be geometric means and the whole progression is geometric...So the common difference becomes a. So, b=a^2, c=a^3, d=a^4, e=a^5 and 1000=a^6

OpenStudy (anonymous):

d= root 10?

OpenStudy (anonymous):

\[\sqrt[3]{10}\]

OpenStudy (anonymous):

did u get it?